OpenStudy (anonymous):

if i had any equation lets say like this 2p + 10 = 24 i dont need it solved but if its plus 10 does that mean you take away 10 from the 24 or add to the 24?

OpenStudy (anonymous):

take away 10 from both sides

OpenStudy (anonymous):

Thats what i thought but i didnt want to get a question like that wrong.

OpenStudy (anonymous):

use additive inverse property of addition. .

jhonyy9 (jhonyy9):

2p+10=24 2p+10-10=24-10 2p=14 -divide both sides by 2 2p/2 =14/2 p=7